Flint Fabricators Inc. machines metal parts for the automotive industry. Under the traditional manufacturing approach, the parts are machined through two processes: milling and finishing. Parts are produced in batch sizes of 60 parts. A part requires 3 minutes in milling and 6 minutes in finishing. The move time between the two operations for a complete batch is 7 minutes.

Under the lean philosophy, the part is produced in a cell that includes both the milling and finishing operations. The operating time is unchanged; however, the batch size is reduced to 6 parts and the move time is eliminated.

Determine the value-added, non-value-added, and total lead times, and the value-added ratio under the traditional and lean manufacturing methods. If required, round percentages to one decimal place.

Traditional Philosophy Lean Manufacturing Philosophy
Value-added time min min
Non-value-added time min min
Total lead time min min
Value-added ratio (as a percent) % %
Answer #1

ANSWER:

Traditional Philosophy

Lean Manufacturing Philosophy

Value-added time

3+6 =9

3+6=9

Non-value-added time

9*(60-1)+7=538

9*(6-1) =45

Total lead time

538+9=547

9+45=54

Value-added ratio (as a percent)

9/547*100 = 1.64%

9/54*100=16.67%

_____________________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU**************